OpenStudy (anonymous):

a wheel has a radius of 25 feet, its rotating at two revolutions per minute. find the linear speed in feet per minute

OpenStudy (jdoe0001):

\(\bf \textit{linear speed} = \cfrac{arc\ length}{time} \implies\cfrac{s}{t}\\\\ \quad \\ \textit{if the radius is 25feet, what's the circumference?}\\ circumference = 2\pi r \implies 2\pi (25) \implies 50\pi\\ \textit{so it does } 50\pi \textit{ feet per in revolution, that is } 2\pi \implies \cfrac{50\pi}{rev}\\\\ \quad \\ \textit{if it does } 50\pi \textit{ feet in 1 revolution, how much would it do in 2 revolutions?}\)